You invest $ 5000 in Acme Inc. on January 1, 2000. Your investment returns 2.75 % compounded monthly. How much money will you have on June 30, 2006?
Answer by mananth(16946)


r = R/100
r = 2.75/100
r = 0.0275 rate per year
Amount A = P(1 + r/n)nt
A = 5,000.00(1 + 0.0275/12)^((12)(6.5))
A = 5,000.00(1 + 0.0023)^(78)
A=$5981